OpenStudy (anonymous):

this is for someone else. Abby's garden has an area of 72 ft^2. Find the length of the garden if the length is twice the width. thanks

OpenStudy (anonymous):

Answer assumes this garden is rectangular, since you describe length and width Let W be width, and L be length then A = LW. Since length is twice width, L = 2W. Substitute for L and A = 2W^2 Since area is 72, A = 72, and now 72 = 2W^2. Divide by 2. 36 = W^2. Square root both sides. 6 = W. Since L = 2W, substitue and find L = 12 ft